### v2

# Search for neutrino emission from blazar $\gamma$-ray flares accounting for possible neutrino time delays

Abstract
We report the results of the search for the high-energy neutrino emission associated with blazar flares, accounting for a possible lag of neutrinos with respect to the electromagnetic emission, either due to the slowness of the proton energy losses in $p\gamma$ collisions and/or proton acceleration. We perform two tests, cross-matching neutrinos with energies $E_{\nu}\gtrsim 100$ TeV from the public catalogue of neutrino alerts IceCat-1 with active galactic nuclei from two source samples based on 1) the MOJAVE database and 2) the CGRaBS catalogue, and utilising Fermi-LAT light curves from the public light curve repository. We scan over a wide range of values of the jet-frame time delay $t^{\prime}_{\mathrm{delay}}$ between the neutrino arrival and the time of the prior major $\gamma$-ray flare and find a pre-trial $\sim 2\sigma$ correlation at $t^{\prime}_{\mathrm{delay}}\sim 10^{3}$ d, which is consistent ($p_{\mathrm{post-trial}}\sim 0.1$) with expectations under the null hypothesis after trial correction.
keywords:
active galactic nuclei, neutrino astronomy
1 Introduction
In the 2010s, the IceCube neutrino observatory discovered astrophysical high-energy ($E_{\nu}\gtrsim 1$ TeV) neutrinos (Aartsen and others, 2013). Around $\sim 5-30$% of their flux was later attributed to the Milky Way galaxy (Abbasi and others, 2023b; Kovalev et al., 2022; Allakhverdyan and others, 2025). Aside from that, besides the evidence of an individual association of the multi-wavelength flare in blazar TXS 0506+056 with a $\gtrsim 200$ TeV neutrino (Aartsen and others, 2018) and evidence of 0.31000.3-100 TeV neutrino emission from active galaxies NGC 1068 (Abbasi and others, 2022), NGC 3079 and NGC 4151 (Neronov et al., 2024), the main contributors to the observed neutrino flux remain unknown (for a recent review, see, e.g., Troitsky, 2021, 2024; Kurahashi et al., 2022).
Active galactic nuclei (AGNs) and their subset with relativistic jets pointing to the Earth at a small angle, blazars, were proposed as neutrino sources almost half a century ago (Berezinsky and Ginzburg, 1981). Still, it was not until recently when some evidence of their detectable contribution to the observed high-energy neutrino flux has been found (Plavin et al., 2020, 2021, 2023, 2024, 2025; Hovatta and others, 2021; Giommi et al., 2020; Kouch and others, 2024; Buson and others, 2022; Buson et al., 2022; Suray and Troitsky, 2023; Albert and others, 2024). Yet, the (in)significance of the blazar-neutrino association depends on neutrino datasets, blazar catalogues and test statistics used in searches for correlations between AGNs and neutrinos (Bellenghi et al., 2023; Abbasi and others, 2023c).
Previous studies, looking for associations between blazars and IceCube high-energy neutrinos, used as proxies either time-averaged fluxes of sources in various electromagnetic bands (from radio to $\gamma$-ray) or their light curves with corresponding values close to the time of the neutrino arrival. However, from the theoretical point of view, the temporal coincidence of the neutrino arrival and the major electromagnetic flare (here, we focus on $\gamma$-ray flares; note that radio flares tend to lag behind $\gamma$-ray flares (Kramarenko et al., 2021) and have longer duration, so we do not discuss them here) might not be warranted for at least two reasons: (i) slow proton acceleration, and/or (ii) slow proton energy losses.
The time required to accelerate protons to energies corresponding to the energies ($\gtrsim 10^{14}$ eV) of neutrinos produced in $p\gamma$ collisions is much larger than the acceleration time of electrons responsible for electromagnetic flares. Given that a neutrino with energy $E_{\nu}$ is typically produced by a proton with energy $E_{p}\sim 20E_{\nu}$, and that the energy of the neutrino produced in a jetted source with Doppler factor $D$ and redshift $z$ in the jet frame is $E_{\nu}^{\prime}=E_{\nu}(1+z)/D$ [^1] Throughout the text, the primed quantities are defined in the rest frame of the jet plasma motion of the corresponding source (shortly, the jet frame)., we obtain $E_{\nu}\sim E_{p}^{\prime}D/(20[1+z])$, i.e. the observed neutrino energy $E_{\nu}$ is roughly equal to the jet-frame parent proton energy $E_{p}^{\prime}$ for typical values of $D\sim 30$, $z\sim 0.5$. For most blazars, their electromagnetic emission during flares can be successfully explained with leptonic models with electrons reaching energies $E^{\prime\max}_{e}\sim 10^{9}-10^{13}$ eV (depending on the blazar class and model assumptions). If the protons are accelerated in the same zone as the electrons via the same mechanism, then, e.g., in the assumption of the linear dependence of the acceleration timescale on the maximum energy, $t_{\mathrm{acc}}^{\prime}\propto E^{\prime\max}$, protons will take one to five orders of magnitude longer time to accelerate to energies required for the neutrino production than electrons to reach energies necessary for explaining the flaring $\gamma$-ray emission.
The characteristic timescale of proton energy losses $t^{\prime}_{p\gamma}$ may be several orders of magnitude larger than the timescale of electron energy losses responsible for the observed electromagnetic flares. During major flares in blazars, injected electrons typically cool within days or less in the observer frame, but protons need a much longer time to interact with photons for effective production of neutrinos if it happens in the energy range of $\sim 100$ TeV near the threshold of the photopion production on optical/ultraviolet/X-ray photon fields.
Whether it is the acceleration timescale $t^{\prime}_{\mathrm{acc}}$ or the $p\gamma$ energy-loss timescale $t^{\prime}_{p\gamma}$ which causes the possible time lag $t^{\prime}_{\mathrm{delay}}$ of the neutrino arrival depends on the conditions in the source, such as the acceleration mechanism, magnetic field, and target photon fields, and it is the slowest process which will determine the neutrino time lag, i.e. $t^{\prime}_{\mathrm{delay}}\sim\max\{t^{\prime}_{\mathrm{acc}},\,t^{\prime}_{p\gamma}\}$. Depending on the model parameters, $\gtrsim 100$ TeV neutrinos can arrive months to years later than the corresponding electromagnetic flare peak (Podlesnyi and Oikonomou, 2025). A study of a possible association of a few lagging neutrinos with another class of sources, namely tidal disruption events, was recently performed by Lu et al. (2025).
If one argues that the proton acceleration timescale is not substantially larger than the electron acceleration timescale, and the $p\gamma$ collisions happen far from the threshold and efficiently, then protons can be accelerated to energies as high as $\sim 100$ PeV and produce neutrinos of corresponding energies (with a negligible time delay with respect to the electromagnetic flare), since the peak energy of the neutrino spectral-energy distribution closely follows the maximum proton energy (Podlesnyi and Oikonomou, 2025; Rodrigues et al., 2025). However, only one neutrino with energy $E_{\nu}\gtrsim$100 PeV has been discovered so far, namely KM3-230213A (KM3NeT Collaboration, 2025). Thus, some of those $\gtrsim 100$ TeV astrophysical neutrinos detected by the IceCube observatory (Abbasi and others, 2023a), if they are produced in a causal connection with major flares in blazars, can be expected to come with some time delay with respect to the flare peaks, either due to the large proton acceleration or $p\gamma$ energy-loss timescale.
The publicly available Fermi-LAT light-curve repository (LCR) (Abdollahi and others, 2023) with regular-binned light curves of all significantly variable Fermi-LAT-detected AGNs spanning more than 15 years allows us to search for association of major $\gamma$-ray flares with lagging behind them IceCube high-energy neutrinos published by the IceCube collaboration in the first alert catalogue IceCat-1 (Abbasi and others, 2023a; IceCube Collaboration, 2023). In this paper, we make the first attempt at this search. In Sect. 2, we present the datasets and catalogues used in our work; in Sect. 3, we present the methods used for searching for associations of lagging neutrinos with major $\gamma$-ray flares in AGNs; in Sect. 4, the results are shown. We discuss implications of our findings and conclude in Sect. 5.
2 Data
2.1 Neutrino alerts from the IceCat-1 catalogue
We use the IceCat-1 catalogue of high-energy neutrino alerts published by the IceCube collaboration (Abbasi and others, 2023a; IceCube Collaboration, 2023). In the following, we use the terms “alert” and “neutrino” interchangeably. We remove eight entries for which there is a flag indicating a coinciding signal in the IceTop detector of cosmic rays, implying a high chance of those alerts being background. This leaves us with $N_{\nu}=348$ neutrino alerts. For each alert $\nu$, the following information of interest is available:
$x_{\nu}=\{s_{\nu},t_{\nu},E_{\nu},\alpha_{\nu},\delta_{\nu},\vec{\epsilon}_{\nu}\},$ (1)
where $s_{\nu}$ is the neutrino signalness reported by the IceCube collaboration in the IceCat-1 catalogue which varies from $0.126$ to $0.997$ with a median value of $0.410$ and represents the probability of the neutrino $\nu$ being of astrophysical origin assuming an $E^{-2.19}_{\nu}$ astrophysical neutrino power-law flux (although this assumption does not hold, e.g., in cases when neutrinos are produced in $p\gamma$ interactions near the threshold, see, e.g., Rodrigues et al., 2024a, b; Kuhlmann and Capel, 2025), $t_{\nu}$ is the moment of the neutrino detection, $E_{\nu}$ is the estimated muon energy, which is a good proxy of the neutrino energy (see Fig. 7 by Abbasi and others, 2023a), $\alpha_{\nu}$ is the right ascension (RA), and $\delta_{\nu}$ is the declination of the most probable sky coordinate of the neutrino arrival, while the vector $\vec{\epsilon}_{\nu}=\{\alpha_{\nu}^{+},\,\alpha_{\nu}^{-},\,\delta_{\nu}^{+},\,\delta_{\nu}^{-}\}$ incorporates information about the uncertainty of the reconstructed direction of the neutrino arrival, where $(\alpha_{\nu}^{+}+\alpha_{\nu}^{-})$ defines the width, and $(\delta_{\nu}^{+}+\delta_{\nu}^{-})$ defines the length of the minimum rectangle encapsulating the neutrino uncertainty arrival contours (see Fig. 1 by Abbasi and others, 2023c).
2.1.1 Set of mock neutrino datasets
Based on the alert dataset $\{x_{\nu}\}$, we create a set of $N_{f}=49\,999$ mock datasets $\{m_{\nu}^{f}\}$, $f=1,\,...,\,N_{f}$, where for each mock alert all the parameters $x_{\nu}$ remained unchanged except for the RA $\alpha_{\nu}^{f}$ and the arrival time $t_{\nu}^{f}$, i.e.
$m_{\nu}^{f}=\{s_{\nu},t_{\nu}^{f},E_{\nu},\alpha_{\nu}^{f},\delta_{\nu},\vec{\epsilon}_{\nu}\}.$ (2)
Since the IceCube detector is located at the South Pole, its sensitivity to astrophysical neutrinos on a long timescale depends only on the energy and declination. Thus, following Plavin et al. (2020), by generating a random uniformly distributed number between $0^{\circ}$ and $360^{\circ}$ for each of the values of $\alpha_{\nu}^{f}$, we obtain a set of mock neutrino datasets reflecting a situation of neutrinos registered from random sky directions (losing a possible connection with a potential class of sources) but keeping the dependence of the IceCube sensitivity on the declination unchanged. For correlation searches involving the time of the neutrino arrival, we generate random values of $t_{\nu}^{f}$ from a uniform distribution $U[t_{\nu}^{\mathrm{min}},t_{\nu}^{\mathrm{max}}]$, where $t_{\nu}^{\mathrm{min}}=55695.064$ (MJD) is the arrival time of the first neutrino in the IceCat-1, and $t_{\nu}^{\mathrm{max}}=60231.917$ (MJD) is the time of the last alert in IceCat-1. Thus, the total time span of IceCat-1 is $T_{\mathrm{\textit{IceCube}}}\approx 4537$ d $\approx 12.4$ yr. Using the Kolmogorov-Smirnov test, we verified, that the original set of RAs of IceCube neutrinos $\{\alpha_{\nu}\}$ and their arrival times $\{t_{\nu}\}$ cannot be distinguished from a set drawn from the corresponding uniform distribution.
2.2 Source lists
2.2.1 MOJAVE-based source list
Since the aim of the paper is to search for blazar-neutrino associations with account for a possible neutrino lagging, which we assume to be universal in the jet frame, to transform the anticipated neutrino arrival delay from the jet frame into the observer frame, the measurements of the blazar’s Doppler factor $D$ and redshift $z$ are required, since $t_{\mathrm{\,delay}}=t^{\prime}_{\mathrm{delay}}(1+z)/D$. As a starting point, we use the sample of $447$ AGNs compiled by Homan et al. (2021), which consists of radio-bright AGNs monitored by the MOJAVE program. All sources in the selected list are located at declinations $30^{\circ}$ and have a minimum 15 GHz correlated flux density $\gtrsim 50$ mJy. All these sources are also present in the Radio Fundamental Catalogue (RFC), version 2025a (Petrov and Kovalev, 2025). We utilize Table 4 in the work of Homan et al. (2021), where the Doppler factors of the blazars are reported, obtained under the assumption that all sources have observed brightness temperature $\propto DT^{\prime}_{b}/(1+z)$, and the intrinsic brightness temperatures $T^{\prime}_{b}$ are determined according to Kovalev and others (2005).
As the next step, we use the 4FGL-DR4 catalogue of Fermi-LAT sources (Abdollahi and others, 2020) available at the mission website [^2] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/14yr_catalog/gll_psc_v35.fit as the file gll_psc_v35.fit. We cross-match 4FGL sources with those presented in the catalogue of Homan et al. (2021).
We use the 4LAC catalogue (Ajello and others, 2020) to obtain the redshift values for the sources in our list of blazars. For those objects which do not have a finite redshift value in the 4LAC catalogue, we use the astroquery.ned module to query the redshift from the NASA/IPAC Extragalactic Database (NASA/IPAC Extragalactic Database, 2019).
Finally, we use the Fermi-LAT LCR (Abdollahi and others, 2023) and download monthly-binned photon flux light curves. We leave only those objects which are present in the LCR, i.e. those with significant variability. The final list of 4FGL AGNs with known redshifts, $\gamma$-ray light curves and Doppler factors reported by Homan et al. (2021) consists of $N_{b}=294$ sources. Among them, according to the classification in the 4FGL catalogue, there are 184 flat-spectrum radio quasars (FSRQs), 93 BL Lacertae-like objects (BLLs), 6 radio galaxies (RDGs), 4 blazars of uncertain type (BCUs), 4 Seyfert galaxies of type 1 (NLSY1s), 2 compact steep-spectrum radio sources (CSSs), and 1 steep-spectrum radio quasar (SSRQ). Since most of the sources are blazars, we use the terms “source”, “AGN”, and “blazar” interchangeably throughout the paper. For blazars for which only lower limits on Doppler factors were reported by Homan et al. (2021) due to unknown redshift, we multiply the lower limits on $D$ by $(1+z)$ with the values of $z$ we found in the NASA/IPAC Extragalactic Database [^3] This is because the observed brightness temperature $\propto DT^{\prime}_{b}/(1+z)$ was assumed by Homan et al. (2021) to correspond to $z=0$ for sources with unknown redshifts, hence the lower limit on $D$ was reported.. The obtained blazar list is denoted as H21+.
2.2.2 CGRaBS-based source list
Homan et al. (2021) obtained Doppler factors under the assumption that the apparent brightness temperature at $15$ GHz depends only on the blazar jet Doppler factor $D$ and the intrinsic (jet-frame) brightness temperature, which for all blazars in the sample is equal to $4\times 10^{10}$ K, as found according to Kovalev and others (2005). This approach can give only approximate values of the Doppler factor because it does not allow for intrinsic source-to-source variations and, besides, may result in incorrect estimations of $D$ when radio images of the jet and the core of the blazar are superimposed (Kovalev et al., 2025).
Thus, it is of interest to search for blazar-neutrino associations using a sample of sources with their Doppler factors obtained with a different method. Recently, Rodrigues et al. (2024b) modelled time-averaged spectral-energy distributions (SEDs) for a sample of 324 blazars, obtaining a fit to the multi-wavelength SEDs with a leptonic and a leptohadronic model, determining values of the Doppler factor $D$ (among other parameters) for each source. The source sample of Rodrigues et al. (2024b) is based on the flux-limited ($F_{8.4\mathrm{\,GHz}}>65$ mJy) CGRaBS catalogue of radio-loud AGNs (Healey et al., 2008) which was cross-matched by Paliya et al. (2017) with sources from various Fermi-LAT catalogues.
We cross-match sources from the list of 324 blazars (Rodrigues et al., 2024b) [^4] Their Doppler factors and redshifts available at https://github.com/xrod/lephad-blazars/blob/main/model_parameters.csv. with the Fermi-LAT LCR (Abdollahi and others, 2023). This results in $N_{b}=295$ sources with known $D$, $z$ and light curves in the LCR. Among them, according to the Fermi-LAT classification (Abdollahi and others, 2020), there are $218$ FSRQs, $72$ BLLs, $2$ BCUs, and $3$ NLSY1s. We denote the compiled source list as R24+.
3 Methods
In this section, we describe our procedure for searching for correlations between blazars in the given blazar list and neutrino alerts from IceCat-1. Our results follow in Sect. 4.
Before doing any statistical test, we perform a spatial association of the neutrinos (both real $\{x_{\nu}\}$ and mock $\{m^{f}_{\nu}\}$) with the sources in the given blazar list (for each of the two tested source lists H21+ and R24+ separately): blazar $b$ is considered to be associated with neutrino $\nu$ if
$\left(\frac{d^{\alpha}_{\nu b}}{\tilde{\alpha}_{\nu}^{\pm}}\right)^{2}+\left(\frac{d_{\nu b}^{\delta}}{\tilde{\delta}_{\nu}^{\pm}}\right)^{2}\leq 1,$ (3)
where
$\tilde{\alpha}_{\nu}^{\pm}=\alpha_{\nu}^{\pm}+\Delta,$ (4)
$\tilde{\delta}_{\nu}^{\pm}=\delta_{\nu}^{\pm}+\Delta,$ (5)
where $d^{\alpha}_{\nu b}$ is the difference between blazar’s $b$ and neutrino’s $\nu$ RAs, $d^{\delta}_{\nu b}$ is the difference between blazar’s and neutrino’s declinations, the sign ++ or - in $\alpha_{\nu}^{\pm}$ and $\delta_{\nu}^{\pm}$ is chosen according to the sign of $d^{\alpha}_{\nu b}$ and $d^{\delta}_{\nu b}$; $\Delta=0.78^{\circ}$ is the assumed systematic angular uncertainty. This value was found (Plavin et al., 2023) to give the optimal correlation strength between the radio-bright blazars and neutrino alerts, and we keep it fixed to avoid statistical penalising for multiple trials. A similar value of $1^{\circ}$ was found (Kouch and others, 2024) to be optimal in the search for spatio-temporal correlations between the multiwavelength light curves of the CGRaBS blazars (Healey et al., 2008) and the IceCube high-energy neutrinos. In a dedicated study with synthesised catalogues and a controlled number of synthetic blazar-neutrino correlations, Kouch et al. (2025b) showed that enlarging the value of $\Delta$ does not lead to an increase in chances of finding a false-positive statistically significant correlation.
Following Eq. (1) in the work by Kouch et al. (2025b), we define for each of the alerts its enlarged angular uncertainty area as
$U_{\nu}=\frac{\pi}{4}\left(\tilde{\alpha}^{+}_{\nu}\tilde{\delta}^{+}_{\nu}+\tilde{\alpha}^{-}_{\nu}\tilde{\delta}^{+}_{\nu}+\tilde{\alpha}^{-}_{\nu}\tilde{\delta}^{-}_{\nu}+\tilde{\alpha}^{+}_{\nu}\tilde{\delta}^{-}_{\nu}\right),$ (6)
which is used in our tests in factor $\rho_{\nu}$ penalising alerts with a large arrival uncertainty.
If neutrinos are produced in the external photon fields, approximate linearity between the $\gamma$ and $\nu$ fluxes may be expected. According to Dermer et al. (2012), the Doppler amplification $\propto D^{p}$ makes the observed neutrino flux boosted with $p=4$ if neutrinos are produced in $p\gamma$ interactions on internal photon fields co-moving with the jet frame. However, $p\geq 5$ is expected if $p\gamma$ interactions are dominated by the photon fields external to the relativistic jet, with especially strong amplifications $p>5$ if $p\gamma$ photopion production occurs near the threshold. The $\gamma$-ray external Compton emission is, in turn, produced with amplification $\propto p^{6}$, making $F_{\nu}\propto F_{\gamma}^{\geq 5/6}\sim F_{\gamma}$, which justifies our assumption of the linear $\gamma-\nu$ flux scaling. Franckowiak and others (2020) found that neutrino-emitting blazar candidates are statistically compatible with hypotheses of both a linear correlation and no correlation between neutrino and gamma-ray energy fluxes.
We expect the jet-frame time delay $t^{\prime}_{\mathrm{delay}}$ to relate to the photopion proton energy-loss timescale $t^{\prime}_{p\gamma}$ and/or proton acceleration timescale $t^{\prime}_{\mathrm{acc}}$. They can vary significantly depending on the proton energy and parameters of photon fields serving as the target for the photopion production process, on the magnetic field, and proton acceleration mechanism. A priori, we do not know the exact value of $t^{\prime}_{\mathrm{delay}}$. Thus, we perform a scan over parameter $t^{\prime}_{\mathrm{delay}}\sim U[t^{\prime}_{\min};t^{\prime}_{\max}]$, $t^{\prime}_{\min}=\Delta_{F}=1$ month corresponding to the bin size $\Delta_{F}$ of the Fermi-LAT light curves. Podlesnyi and Oikonomou (2025), in modelling the brightest Fermi-LAT blazar flare of 3C 454.3, obtained that close to the threshold of the photopion production around $E_{\nu}\sim 100$ TeV, $t^{\prime}_{\mathrm{delay}}\approx t^{\prime}_{p\gamma}\approx t^{\prime}_{\mathrm{acc}}\sim 10^{9}$ s (see Fig. 12 by Podlesnyi and Oikonomou, 2025). Thus, we adopt $t^{\prime}_{\max}=3\times 10^{9}$ s $3.47\times 10^{4}$ d being agnostic of the exact mechanism of the possible $\nu$ delay and no particular preference to any delay between $t^{\prime}_{\min}$ and $t^{\prime}_{\max}$.
For each alert $\nu$, which has at least one source from the given source list in its uncertainty region (hereafter “associated alert”), we assign a weight
$w_{\nu}(t^{\prime}_{\mathrm{delay}})=\sum_{a}\left[s_{\nu}\rho_{\nu}G_{\nu a}(t^{\gamma\max}_{a},t_{\nu};t^{\prime}_{\mathrm{\,delay}},D_{a},z_{a})F^{\gamma}_{a}(t^{\gamma\max}_{a})\right],$ (7)
where the sum $\sum_{a}$ runs over all sources within the uncertainty region of neutrino $\nu$ (as was pointed out by Kouch et al. (2025a), a single neutrino cannot come from more than one source, but having several blazars in its angular uncertainty region increases the probability of this neutrino coming from a blazar), $s_{\nu}$ is the alert signalness, $\rho_{\nu}$ is the factor penalising alerts with too large uncertainties. Following recommendations of Kouch et al. (2025b), we use a top-hat spatial factor penalising alerts with too large angular uncertainties:
$\rho_{\nu}=\begin{cases}1,\mathrm{\,if\,}U_{\nu}\leq\tilde{U}_{\nu}\\ \displaystyle{\left(\frac{\tilde{U}_{\nu}}{U_{\nu}}\right)},\mathrm{\,if\,}U_{\nu}>\tilde{U}_{\nu},\end{cases}$ (8)
where $\tilde{U}_{\nu}=15.6$ sq. deg. is the median angular uncertainty (see Eq. 6) of IceCat-1 alerts (after the extension by $\Delta=0.78^{\circ}$).
The key point of the current paper is the search for neutrino-blazar correlations taking into account the possibility of neutrino emission lagging behind the $\gamma$-ray flare due to slowness of the $p\gamma$ photopion production and/or proton acceleration as was suggested by Podlesnyi and Oikonomou (2025). Given that even during the brightest blazar flares, the IceCube neutrino yield is of the order of $10^{-3}-10^{-2}$ muon neutrinos at energies $\gtrsim 100$ TeV (Oikonomou et al., 2019; Kreter et al., 2020; Robinson and Boettcher, 2024) and during quieter activity periods even lower (assuming proportionality between $\gamma$-ray and $\nu$ fluxes), we assume that if the neutrino is indeed associated with the blazar, it must have come in a causal connection with the major flare. Thus, the last term in Eq. (7) denotes the global maximum of the Fermi-LAT light curve of blazar $a$ detected at $t^{\gamma\max}_{a}$ among the times prior to the associated neutrino arrival ($\Delta_{F}=1$ month to account for the size of the time bin and still allow for instantaneous associations) normalized by $1.66\times 10^{-5}$ cm-2 s-1 (the maximum value of monthly-averaged fluxes among the studied blazars within the whole Fermi-LAT observation history observed from FSRQ 3C 454.3 during the brightest Fermi-LAT blazar flare in November 2010 (Fermi-LAT Collaboration, 2011; Nalewajko, 2013; Podlesnyi and Oikonomou, 2025)). As for the temporal term $G_{\nu a}$ accounting for the neutrino delay, we define it linking to the time $t^{\gamma\max}_{a}$ of the global maximum of the associated blazar light curve prior to the neutrino arrival as follows
[…]
where the delay for source $a$ in the observer’s frame is given by
$t_{a\mathrm{\,delay}}(t^{\prime}_{\mathrm{delay}},D_{a},z_{a})=\max\left\{\Delta_{F},\frac{1+z_{a}}{D_{a}}t^{\prime}_{\mathrm{delay}}\right\}.$ (10)
Figure 1: An example of the graph of the temporal part $G_{\nu a}$ of weights defined in Eq. (9) as a function of $t^{\gamma\max}_{a}$ for a fixed $t_{\nu}$ (shown as the red solid vertical line) and a set of fixed $t_{a\mathrm{\,delay}}$ (see the legend).
In Eq. (9) in the first row we, for the sake of completeness, set $G_{\nu a}$ as zero for $t^{\gamma\max}_{a}>t_{\nu}+\Delta_{F}$, albeit technically $G_{\nu a}$ is not defined for these times because by definition $t^{\gamma\max}_{a}$ is the time of the maximum of the Fermi-LAT light curve of blazar $a$ only among times $\leq t_{\nu}+\Delta_{F}$. The second row in Eq. (9) assigns equal weights to all values of $t^{\gamma\max}_{a}$ if they preceed $t_{\nu}$ by $t_{a\mathrm{\,delay}}$ with an absolute margin of $\Delta_{F}$ to account for the finite size of the Fermi-LAT time bin. The third row handles all other cases assigning the Gaussian penalty with the width equal to the anticipated neutrino delay $t_{a\mathrm{\,delay}}$ in the Earth frame. The illustration of the graph of $G_{\nu a}$ as a function of $t^{\gamma\max}_{a}$ for a fixed $t_{\nu}$ and $t_{a\mathrm{\,delay}}$ is given in Fig. 1.
As was demonstrated by Kouch et al. (2025b), the most sensitive global test statistic (TS) for the whole dataset is just the sum of individual weights:
$\mathrm{TS}(t^{\prime}_{\mathrm{delay}})=\sum\limits_{\nu}w_{\nu}(t^{\prime}_{\mathrm{delay}}),$ (11)
and TS is a function of the jet-frame neutrino time delay $t^{\prime}_{\mathrm{delay}}$, assumed to be universal among all sources.
We compare the obtained $\mathrm{TS}$ value for the real dataset $\{x_{\nu}\}$ against the distribution of values $\mathrm{TS}_{f}$ obtained for the set of mock datasets $\{m_{\nu}^{f}\}$ with values of right ascensions $\alpha_{\nu}^{f}$ and arrival times $t_{\nu}^{f}$ randomly generated as described in Sect. 2.1.1. Then, for each correlation search (test) performed, we, following Plavin et al. (2020); Davison and Hinkley (1997), obtain the p-value
$p(t^{\prime}_{\mathrm{delay}})=\frac{\#\{\mathrm{TS}_{f}\geq\mathrm{TS}\}_{f=1}^{N_{f}}+1}{N_{f}+1},$ (12)
i.e. $p(t^{\prime}_{\mathrm{delay}})$ is the fraction of times the test statistic for a mock dataset $\mathrm{TS}_{f}^{i}$ was larger than or equal to the TS value for the real dataset $\{x_{\nu}\}$ at the given value of the $\nu$ time delay $t^{\prime}_{\mathrm{delay}}$.
4 Results
The results of the tests performed for H21+ and R24+ source lists are shown in Fig. 2. For the H21+ source list, the global minimum of p-value is $0.037$ reached at the jet-frame delay $\tilde{t}^{\prime}_{\mathrm{delay}}=7.7\times 10^{3}$ d. For the R24+ source list, the minimum of $p(\hat{t}^{\prime}_{\mathrm{delay}})=0.026$ is reached at $\hat{t}^{\prime}_{\mathrm{delay}}=1.9\times 10^{3}$ d. Tables 1 and 2 present the top ten associations with the highest values of $\mathrm{TS}_{\nu}$ for H21+ and R24+ accordingly [^5] For the H21+ source list, we ran an additional test with an extra factor $F^{X}_{a}$ in Eq. (7), where $F^{X}_{a}$ is the relative radio flux of sources in the X radio band (between 8 GHz and 12 GHz). This test was motivated by Plavin et al. (2020, 2023), that reported evidence of the correlation between the IceCube neutrino alerts and radio fluxes in the cores of radio sources at the frequency of $8$ GHz. From the RFC, we used the column with the “median flux density at X-band at baseline projection lengths shorter than 1000 km” (Petrov and Kovalev, 2025) and normalised it by the maximum value among the sources in the H21+ source list. This test did not result in a significant correlation with a p-value curve similar to the one in Fig. 2..
Figure 2: $p$-value (see Eq. 12) as a function of the jet-frame time delay assumed to be universal for all AGNs for the test with the sources from H21+ (dotted curve) and sources from R24+ (dashed-dotted curve).
In Fig. 2, one can see that the p-value curve of R24+ almost reaches the $2\sigma$ level (one-tailed). The trial correction (for multiple tested values of $t^{\prime}_{\mathrm{delay}}$) performed according to the prescriptions of Plavin et al. (2020) shows, however, that the minima of the pre-trial p-values as deep as in Fig. 2 occur by chance in $p_{\mathrm{post-trial\,H21+}}=14$% ($1.06\sigma$) and $p_{\mathrm{post-trial\,R24+}}=11$% ($1.21\sigma$) among mock datasets for the tests with H21+ and R24+ source lists respectively. Thus, none of the obtained p-values show evidence of a significant association between IceCat-1 neutrinos and AGNs from the considered source lists.
We note that the optimal delay $\tilde{t}^{\prime}_{\mathrm{delay}}=7.7\times 10^{3}$ d for H21+ is by $4$ times greater than $\hat{t}^{\prime}_{\mathrm{delay}}=1.9\times 10^{3}$ d for R24+. This is driven by the substantial difference in the Doppler factors (compare Tables 1 and 2) of 3C 454.3 and 3C 279, that contribute 5353% of the total TS for the former and 8080% of the total TS for the latter: Homan et al. (2021) estimates the Doppler factor of 3C 279 by a factor of five larger than the estimate of Rodrigues et al. (2024b), and the Doppler factor of 3C 454.3 by a factor of three larger than that found by Rodrigues et al. (2024b).
We show examples of two blazar light curves for associated alerts with the two highest values of $\mathrm{TS}_{\nu}$ in both of the two tests performed in Fig. 3.
The most significant association is driven by the alert IC120523B [^6] In the original publication of Abbasi and others (2023c), this alert had a name IC120523A, which was also given to another alert in the catalogue; in the updated version of IceCat-1 (IceCube Collaboration, 2023), the alert has the name IC120523B. associated with the brightest Fermi-LAT blazar flare of the FSRQ 3C 454.3, which happened in November 2010 (Fermi-LAT Collaboration, 2011; Nalewajko, 2013; Podlesnyi and Oikonomou, 2025). Despite large angular uncertainty of $47.2$ sq. deg., the exceptional brightness of 3C 454.3 during that flare and that the neutrino arrival delay in the Earth frame is not far from the anticipated $(t^{\prime}_{\mathrm{delay}}(1+z)/D)$ for the both tested source lists, provides the highest value of $\mathrm{TS}_{\nu}$ in both Tables 1 and 2. 3C 454.3 is well within the original IceCube angular uncertainty contour lying in $0.73^{\circ}$ from the best-fit arrival direction of IC120523B, and this association does not require its extension by $\Delta$.
3C 279 is the second brightest FSRQ in the Fermi-LAT 4FGL catalogue (Abdollahi and others, 2020). The neutrino IC150926A arrives around $100$ d (mind the finite size of one month of the Fermi-LAT light curves) after the bright but short multiwavelength flare of 3C 279 happened around MJD 57 200 detected as well in X-rays and optical photons (Paliya, 2015; Pittori and others, 2018; Singh et al., 2020). The peak of the flare is in perfect agreement with the anticipated neutrino lagging by $(t^{\prime}_{\mathrm{delay}}(1+z)/D)$ for both H21+ and R24+ source lists. As we mentioned above, the four-times difference in estimates of Doppler factors by Homan et al. (2021) and Robinson and Boettcher (2024) seems to drive the difference in the optimal delays providing the minima of p-values in Fig. 2. The combination of the source brightness and the prior $\gamma$-flux global maximum happening at the anticipated time, together with a relatively small angular uncertainty of $8.8$ sq. deg., result in the second highest value of $\mathrm{TS}_{\nu}$ in both Tables 1 and 2. This association requires the extension of the original IceCube angular uncertainty contour by $\Delta=0.78^{\circ}$.
Alert Association $s_{\nu}$ $E_{\nu}$ [TeV] $U_{\nu}$ $^{{\circ}^{2}}$] $\mathrm{\mathrm{TS}_{\nu}}$ $D$ $z$
IC120523B 3C 454.3 0.490 168 47.2 0.1287 45.3 0.86
IC150926A 3C 279 $\dagger$ 0.296 216 8.8 0.0617 140.2 0.54
IC170922A PKS 0502+049 $\dagger$ 0.631 264 7.1 0.0432 24.6 0.95
TXS 0506+056 1.8 0.34
IC221223A B2 2308+34 $\dagger$ 0.795 353 7.5 0.0181 22.8 1.82
IC180608A PKS 0440-00 $\dagger$ 0.396 158 11.3 0.0168 3.5 0.45
IC160727A 4C +14.23 $\dagger$ 0.296 105 14.9 0.0082 13.9 1.04
IC181023B TXS 0518+211 $\dagger$ 0.427 136 11.5 0.0071 2.7 0.11
IC131014A MG1 J021114+1051 $\dagger$ 0.665 293 6.3 0.0071 6.3 0.20
IC190410A PKS 2029+121 0.280 105 44.6 0.0050 28.3 1.22
PKS 2032+107 9.7 0.60
IC120515A OP 313 0.613 194 13.3 0.0039 24.2 1.00
Table 1: Top ten neutrino alerts with the highest values of their $\mathrm{TS}_{\nu}$ in the test with the source list $\{\mathrm{H}2+\}$ corresponding to the local minimum of $p(\tilde{t}^{\prime}_{\mathrm{delay}})=0.037$ reached at $\tilde{t}^{\prime}_{\mathrm{delay}}=7.7\times 10^{3}$ d. The sum $\mathrm{TS}$ for the whole dataset is $0.3577$. The symbol $\dagger$ indicates associations requiring the extension of the original IceCube errors by $\Delta=0.78^{\circ}$.
Alert Association $s_{\nu}$ $E_{\nu}$ [TeV] $U_{\nu}$ $^{{\circ}^{2}}$] $\mathrm{\mathrm{TS}_{\nu}}$ $D$ $z$
IC120523B 3C 454.3 0.490 168 47.2 0.0787 14.7 0.86
IC150926A 3C 279 $\dagger$ 0.296 216 8.8 0.0617 27.8 0.54
IC160814A PKS 1313-333 $\dagger$ 0.607 263 26.3 0.0117 19.3 1.21
IC211208A PKS 0735+17 $\dagger$ 0.502 171 25.1 0.0036 5.8 0.42
IC150812B PKS 2145+06 $\dagger$ 0.831 508 6.5 0.0025 4.0 1.00
IC140927A PKS 0336-01 $\dagger$ 0.481 182 37.2 0.0019 26.8 0.85
IC170427A RX J0011.5+0058 $\dagger$ 0.383 155 40.3 0.0018 17.3 1.49
S3 0013-00 15.0 1.58
IC120605A OL 318 $\dagger$ 0.385 107 21.1 0.0016 19.9 1.41
B2 1015+35B $\dagger$ 3.6 1.23
IC200614A B2 0202+31 0.415 115 65.1 0.0015 20.5 1.47
IC220225A PKS 0215+015 $\dagger$ 0.378 154 27.2 0.0015 14.6 1.72
Table 2: Top ten neutrino alerts with the highest values of their $\mathrm{TS}_{\nu}$ in the test with the source list R24+ corresponding to the global minimum of $p(\hat{t}^{\prime}_{\mathrm{delay}})=0.026$ reached at $\hat{t}^{\prime}_{\mathrm{delay}}=1.9\times 10^{3}$ d. The sum $\mathrm{TS}$ for the whole dataset is $0.1734$. The symbol $\dagger$ indicates associations requiring the extension of the original IceCube errors by $\Delta=0.78^{\circ}$. Note the change of values for $D$ w.r.t. Table 1 and that TXS 0506+056 is absent in R24+.
Figure 3: Light curves of 3C 454.3 (upper panel), and 3C 279 (lower panel), which are among the most significant associations between IceCat-1 neutrinos and AGNs in both the tests (see Tables 1 and 2). The connected markers with error bars show monthly-binned Fermi-LAT photon fluxes in the energy range 100 MeV–100 GeV from the public repository (Abdollahi and others, 2023) (ordinate axis). Dotted vertical lines indicate the time $t_{\nu}$ of the neutrino arrival. Time intervals between the alert and the global maximum of the light curve prior to $t_{\nu}$ are shown as double-pointed arrows. Vertical upper dashed risks indicate the anticipated position of the maximum with the Doppler factor from R24+ and $\hat{t}^{\prime}_{\mathrm{delay}}=1.9\times 10^{3}$ d, and vertical lower dashed-dotted risks — with the Doppler factor from H21+ and $\tilde{t}^{\prime}_{\mathrm{delay}}=7.7\times 10^{3}$ d. The width of the Gaussian temporal factor $G_{\nu a}$ in weights is equal to the distance between the vertical risks and the line of $t_{\nu}$ (see Eq. 9).
5 Discussion and conclusion
The two statistical tests weighing blazar-associated neutrino alerts accounting for the time lag of the neutrino arrival due to slowness of the proton energy losses in $p\gamma$ interactions and/or proton acceleration resulted in no evidence of a significant correlation between IceCat-1 alerts and the AGNs from H21+ or R24+ source lists. The low statistical significance of these associations $(\lesssim 2\sigma)$, given the small size, is in line with predictions of Liodakis et al. (2022) and implies (Abbasi and others, 2023c; Kouch et al., 2025b, a) that only a small fraction of AGNs can be associated with IceCube alerts.
A set of non-mutually-exclusive reasons can explain the null results of our search:
  • Blazars are inefficient sources of $\sim 100$ TeV neutrinos. Indeed, in our recent work (Podlesnyi and Oikonomou, 2025), we estimated from an extrapolation of a leptohadronic model for the brightest blazar flare of 3C 454.3 to the whole population of Fermi-LAT FSRQs that, on long-term average, they produce $\sim\left(\rho_{p/e}/130\right)\times 0.5$% of IceCube neutrinos at energies greater than $100$ TeV, where $\rho_{p/e}$ is the ratio of the proton and electron energy densities, which is a proxy of the baryon loading factor (the ratio of the proton and $\gamma$-ray luminosities). Taking this rough estimate at face value, given $N_{\nu}=348$ IceCat-1 alerts with their median signalness of $0.41$, we obtain $0.71$ expected neutrinos from Fermi-LAT FSRQs, which naively implies $\mathcal{O}(1)$ AGN-neutrino associations, assuming that other types of Fermi-LAT AGNs contribute to the neutrino flux similarly.
  • Blazars are efficient neutrino sources at energies $\gtrsim 10$ PeV. If blazars are efficient accelerators of protons [^7] By this we mean the case when protons are accelerated close to the Hillas limit (Hillas, 1984), when the ratio of the proton acceleration timescale to the gyroperiod at the maximum energy $\xi_{\mathrm{acc}}\lesssim 10^{3}$., then the peak of the neutrino spectral-energy distributions will be around $E_{\nu}\gtrsim 100$ PeV because the protons of similar energies in the jet frame reach the $p\gamma$ reaction threshold and begin efficiently interacting with the abundant infrared photons (see, e.g., Fig. 12 by Podlesnyi and Oikonomou, 2025). Rodrigues et al. (2025) also shows that this allows to decrease the proton power below the Eddington limit alleviating the issue with super-Eddington proton luminosities often required to explain $\sim 100$ TeV neutrinos.
  • Our initial assumption of the universality of the neutrino time delay $t^{\prime}_{\mathrm{delay}}$ in the blazar jet frame may not be valid. Indeed, the conditions in different sources (e.g., acceleration mechanisms and efficiencies, magnetic and photon fields etc.) can vary too much from source to source and the universal value of $t^{\prime}_{\mathrm{delay}}$ may not exist.
  • Even if $t^{\prime}_{\mathrm{delay}}$ is universal for all neutrino sources, its transformation from the jet frame to the observer frame as $t^{\prime}_{\mathrm{delay}}(1+z)/D$ is significantly affected by the value of $D$. Indeed, there are $174$ AGNs which are present both in the H21+ and R24+ source lists, and the mean ratio of values from H21+ and R24+ is $\langle D^{\mathrm{H}21+}/D^{\mathrm{R}24+}\rangle=1.8$ and the standard deviation of $\sigma(D^{\mathrm{H}21+}/D^{\mathrm{R}24+})=2.1$ indicating a rather large discrepancy between the values obtained with the two different methods. The differences in estimates of Doppler factors in blazar jets with different methods are well known (Liodakis et al., 2017), and even within the same method, at different observational epochs, the same source may have various values of $D$ (Jorstad and others, 2013). Our adoption of the Gaussian weight accounting for the delay was meant to allow some room for uncertainty in $D$ but if it is as large as $\sim 100$%, this can hinder attempts to reveal a significant correlation between major $\gamma$-ray flares and arrivals of IceCube neutrinos.
  • Proton acceleration and/or energy dissipation are decoupled from those of electrons. As Capel et al. (2022) pointed out, in many leptohadronic models, the hadronic contribution is not needed to describe the observational electromagnetic data well (Keivani and others, 2018; Ansoldi and others, 2018; Murase et al., 2018; Sahakyan, 2018; Gao et al., 2019; Oikonomou et al., 2019; Cerruti et al., 2019; Reimer et al., 2019; Rodrigues et al., 2019; Petropoulou and others, 2020; Oikonomou et al., 2021). This implies that potential neutrino emission even if some hadrons are present in the jet might be not causally connected to the leptonic emission because, e.g., protons are accelerated in a different region of the jet and/or most efficiently dissipate in zones different from those of electrons. This, in turn, leads to (i) neutrino flux not necessarily related to the electromagnetic flux of the source; (ii) no characteristic time delay $t^{\prime}_{\mathrm{delay}}$ between the neutrino production and electromagnetic flares.
Plavin et al. (2020); Hovatta and others (2021); Kouch and others (2024) found an indication of the association of IceCube high-energy neutrinos with radio flares (see, however, recent work by Abbasi and others (2025a), where such a correlation was not found), which are observed, on average, to lag behind $\gamma$-ray flares by several months in the Earth frame (Kramarenko et al., 2021) as expected in some models (Boula and Mastichiadis, 2022; Tramacere et al., 2022; Zacharias, 2023). Future studies utilising both $\gamma$-ray and radio light curves may increase the sensitivity of searches for neutrino-blazar associations. The releases of the new CAZ catalogue with $858$ Doppler factors (Kouch and others, 2025) and updated IceCat-2 catalogue of IceCube alerts (Abbasi and others, 2025b) will significantly enlarge both the sizes of the ensembles of sources and neutrino alerts allowing to improve the sensitivity of the search for potentially delayed neutrinos.
Other projected or constructed neutrino telescopes, such as the Pacific Ocean Neutrino Experiment (Agostini and others, 2020), the Baikal Gigaton Volume Detector (Avrorin and others, 2011), and the Cubic Kilometre Neutrino Telescope (Adrian-Martinez and others, 2016), Tropical Deep-sea Neutrino Telescope (Ye and others, 2023), IceCube-Gen2 (Aartsen and others, 2021; Ishihara and others, 2023), and the Radio Neutrino Observatory in Greenland (Aguilar and others, 2021) will help to provide the definitive answer to the question about the total contribution of the AGNs to the astrophysical neutrino flux and a role of the possible neutrino delay with respect to the electromagnetic emission.
In conclusion, the two cross-matching tests between IceCat-1 neutrino alerts and AGNs from the MOJAVE-based source sample H21+ (Homan et al., 2021) and CGRaBS-based source sample R24+ (Rodrigues et al., 2024b) with scanning over the jet-frame time delay $t^{\prime}_{\mathrm{delay}}$ between the neutrino arrival time in IceCube and prior major $\gamma$-ray flares resulted in no significant evidence of the correlation.
Acknowledgments
We thank Nikita Vasiliev for multiple discussions during his earlier unpublished work on the student project on searching for spatio-temporal correlations between AGN $\gamma$-ray emission and IceCube neutrinos. We thank Sergey Troitsky, Yuri Kovalev, and Timur Dzhatdoev for helpful discussions.
This research has made use of the Fermi-LAT (Atwood and others, 2009; Ajello and others, 2020; Abdollahi and others, 2020, 2023) data provided by NASA Goddard Space Flight Center, and of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of data from the MOJAVE database, which is maintained by the MOJAVE team (Lister et al., 2018), and the data from the All-Sky Survey of Gamma-Ray Blazar Candidates CGRaBS (Healey et al., 2008). This work has made use of the following software packages: Python (Van Rossum and Drake, 2009), AstroPy (Astropy Collaboration, 2013; Astropy Collaboration and Astropy Contributors, 2018; Astropy Collaboration, 2022), Jupyter (Perez and Granger, 2007; Kluyver et al., 2016), matplotlib (Hunter, 2007), NumPy (Harris and others, 2020), SciPy (Virtanen and others, 2020; Pauli Virtanen et al., 2021), astroquery (Astroquery collaboration and a subset of the astropy collaboration, 2019; Adam Ginsburg et al., 2024), pandas (pandas development team, 2020; McKinney, 2010), snakemake (Mölder et al., 2021), pyLCR (Abdollahi and others, 2023) [^8] https://github.com/dankocevski/pyLCR, and tqdm (Casper da Costa-Luis et al., 2023). This research has made use of NASA’s Astrophysics Data System, INSPIRE-HEP, and Pathfinder (Kartheik G. Iyer et al., 2024). Software citation information aggregated using the Software Citation Station (Wagg and Broekgaarden, 2024b, a) [^9] https://www.tomwagg.com/software-citation-station/.
Data availability
The compiled source lists and all the codes reproducing the results presented in the paper will be published on Zenodo upon acceptance of the manuscript.